OpenStudy (anonymous):

Rectangle A has an area of 16 − x2. Rectangle B has an area of x2 + 2x − 24. In simplest form, what is the ratio of the area of Rectangle A to the area of Rectangle B? Show your work.

OpenStudy (pfenn1):

ratio of areas of A to B\[\frac{16-x^2}{x^2+2x-24}\]I think this can be simplified by factoring the numerator and denominator

OpenStudy (anonymous):

How would I simplify that?

OpenStudy (anonymous):

oh nevermind lol

OpenStudy (anonymous):

@pfenn1 is (4-x)^2 the factorization of 16-x^2?

OpenStudy (pfenn1):

No. \[(4-x)^2=16-8x+x^2\]\[16-x^2=(4-x)(4+x)\]This is a good thing to memorize. When ever you see a expression in the form of (a^2 - b^2), the factors are always (a - b)(a + b)

OpenStudy (anonymous):

man they really reach to make these algebra questions into "word problems" don't they?